OpenStudy (anonymous):

A zoo train ride costs $5 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who took the ride was 17, and the total money collected was $65. What was the number of children and the number of adults who took the train ride that day, and which pair of equations can be solved to find the numbers?

OpenStudy (anonymous):

5 children and 12 adults Equation 1: a + c = 17 Equation 2: 5a − c = 65 12 children and 5 adults Equation 1: a + c = 17 Equation 2: 5a + c = 65 12 children and 5 adults Equation 1: a + c = 17 Equation 2: 5a − c = 65 5 children and 12 adults Equation 1: a + c = 17 Equation 2: 5a + c = 65

OpenStudy (anonymous):

I think b

OpenStudy (anonymous):

so is it b or d

OpenStudy (0_0youscareme):

D.

OpenStudy (anonymous):

Line A is represented by the equation given below: x + y = 4 What is most likely the equation for line B, so that the set of equations has infinitely many solutions? 4x + 4y = 4 4x + y = 4 2x + 2y = 8 x + y = 8